# Joint Pedestrian Trajectory Prediction through Posterior Sampling

### v1

@@ Abstract
Joint pedestrian trajectory prediction has long grappled with the inherent unpredictability of human behaviors. Recent investigations employing variants of conditional diffusion models in trajectory prediction have exhibited notable success. Nevertheless, the heavy dependence on accurate historical data results in their vulnerability to noise disturbances and data incompleteness. To improve the robustness and reliability, we introduce the Guided Full Trajectory Diffuser (GFTD), a novel diffusion model framework that captures the joint full (historical and future) trajectory distribution. By learning from the full trajectory, GFTD can recover the noisy and missing data, hence improving the robustness. In addition, GFTD can adapt to data imperfections without additional training requirements, leveraging posterior sampling for reliable prediction and controllable generation. Our approach not only simplifies the prediction process but also enhances generalizability in scenarios with noise and incomplete inputs. Through rigorous experimental evaluation, GFTD exhibits superior performance in both trajectory prediction and controllable generation. See these results at https://sites.google.com/andrew.cmu.edu/posterior-sampling-prediction.
@@ I INTRODUCTION
Pedestrian trajectory prediction is crucial for human-robot interaction systems such as autonomous driving, etc. The goal is to predict future paths based on previous movements and the environmental context of pedestrians. By accurately predicting pedestrian trajectories, autonomous systems can plan their actions accordingly, ensuring safe and efficient navigation in dynamic environments. However, the unpredictable and complicated nature of human behaviors makes this task challenging, especially with multiple agents involved where their interactions need to be considered.
@@
To simplify the problem, early works [1, 2, 3, 4, 5] focused on marginal pedestrian trajectory prediction, which forecasts the path for each pedestrian independently. Such approaches require a downstream planning module to perform safety checks for every combination of the individual predictions. Even so, some combinations could still produce unrealistic self-collisions and lead to failure in challenging scenarios. As a result, joint pedestrian trajectory prediction, which predicts consistent trajectories for all agents together, has gained attention in the community. A joint metrics term is introduced by [6] into supervision loss, transforming the marginal predictor into a joint predictor. However, joint pedestrian trajectory prediction remains challenging since existing approaches heavily rely on accurate and complete historical data to incorporate temporal and social dynamics. This dependence results in their vulnerability to interference from noisy disturbances and missing data, significantly threatening their effectiveness in real-world applications, such as sensors under adverse weather conditions.
@@
Figure 1: Illustration of existing trajectory prediction framework and our guided full trajectory diffuser framework. (a) Prior multi-agent trajectory prediction methods directly generate entire future trajectories through supervised learning. (b) Previous multi-agent trajectory prediction methods utilizing diffusion models generate future trajectories step-by-step during the denoising process. (c) Our method for multi-agent trajectory prediction incorporates history guidance within the diffusion framework. It predicts entire trajectories and supports additional capabilities in a single model, including Noise-Robust Trajectory Prediction, Incomplete Data Prediction, and Controllable Trajectory Generation.
@@
To deal with the noises, previous efforts involved augmenting data with predefined noise and training models on noisy datasets [7], and it can be further improved by adversarial training procedures [8]. In addition, historical trajectories from sensors could be incomplete. Previous studies  [9, 10, 11, 12] proposed to reconstruct missing data and predict future trajectories during the training phase, requiring a specifically designed model and well-established training strategy. Such approaches optimize per-problem functions, lacking adaptability to diverse tasks across various contexts.
@@
Draw inspiration from the diffusion model with its remarkable capability of capturing the complicated distribution, we propose a unified framework, named, Guided Full Trajectory Diffuser (GFTD) for the joint pedestrian trajectory prediction to better handle the disturbances and incompleteness. GFTD represents the entire trajectory distribution, both historical and future, with one diffusion model. We formulate trajectory prediction and controllable generation as inverse problems and solve them through posterior sampling techniques. Specifically, we sample in-distribution full trajectories based on historical trajectories and priors such as physical constraints and behavioral intentions. Without the necessity for explicit training to handle noisy and/or incomplete inputs, GFTD enables robust prediction and controllable generation—both achievable during the inference time. In a nutshell, our proposed framework streamlines the training process and offers adaptability to various scenarios at inference, providing a solution that can address all challenges without extra training requirements.
@@ II-A Pedestrian Trajectory Prediction
Pedestrian trajectory prediction is crucial for human-robot interaction systems, such as autonomous driving. This task involves forecasting the future movement paths of pedestrians, given their historical movements and the environment. However, it is challenging to predict their motion because of the diverse and unpredictable nature of human behaviors. To address this problem, two mainstream paradigms have been developed. The supervised learning approach [5] aims to minimize the differences between the ground truth trajectories and the predictions using L2 loss, etc. On the other hand, the generative learning approach [2] formulates the prediction as a task of generating conditional distributions of the future trajectory based on past trajectories. Among various generative models, the diffusion model [13] [14] has shown exceptional performance in pedestrian trajectory prediction. Through a conditional reverse diffusion process, diffusion model generates future trajectory distributions from a standard Gaussian distribution, which captures accurate and diverse predictions of future trajectories.
II-B Marginal and Joint Distribution
Recently, joint pedestrian trajectory prediction has gained significant attention. Traditional methods focus on marginal pedestrian trajectory prediction which treat each pedestrian independently, ignoring potential interactions and the impact on future paths. Instead, joint pedestrian trajectory prediction takes these interactions into consideration, aiming for more consistent and interconnected predictions. [6] incorporates joint metrics as the training objective, transforming the marginal trajectory predictor into a joint trajectory predictor. In this paper, we propose Guided Full Trajectory Diffuser (GTFD) for joint pedestrian trajectory prediction, leveraging the strengths of diffusion models in generating accurate and diverse future trajectory distributions while considering the interactions between pedestrians.
@@ III Methods
In this section, we undertake a comprehensive exposition of our novel multi-agent trajectory prediction framework, conceptualizing the prediction task through the lens of spatial-temporal inpainting. We will first introduce the problem formulation and preliminaries to diffusion models. Subsequently, we explain how such transformation is made possible and provide a detailed exposition of our novel prediction model within our framework.
@@ III-A Preliminaries on Diffusion models
According to Itô SDE, the forward data noising process is defined as the following form:
$dx=f(x,t)dt+g(t)dw,$ (1)
where $w$ is the standard Wiener process, $f$ is the drift coefficient, and $g$ is the diffusion coefficient.
@@
$dx=[f(x,t)-g^{2}(t)\nabla_{x}{\log{p_{t}}(x)}]dt+g(t)dw,$ (2)
where $p_{t}(x)$ is the probability density at diffusion step $t$, and $\nabla_{x}{\log{p_{t}}(x)}$ is the time-dependent score function, which is approximated by training a neural network $s_{\theta}(x_{t},t)$ with score matching. Here, $x_{t}$ denotes the intermediate noisy data $x$ at denoise time $t$. The training objective is defined as
$\mathbb{E}_{t}\{\lambda(t)\mathbb{E}_{x_{0}}\mathbb{E}_{x_{t}|x_{0}}[\left\|s_{\theta}(x_{t},t)-\nabla_{x_{t}}{\log{p}(x_{t}|x_{0})}\right\|_{2}^{2}]\},$ (3)
where $\lambda(t)$ is a weighting function, and $t$ is uniformly sampled over $[0,T]$.
In this paper, we follow the implementation introduced by Denoising Diffusion Probabilistic Models (DDPM) [23]. DDPM gradually add noise to the orginal data $x$ according to a variance schedule $\beta_{1}$, $\beta_{2}$,…,$\beta_{T}$:
$x_{t}=\sqrt{1-\beta_{t}}x_{t-1}+\sqrt{\beta}_{t}\epsilon,\quad t\in\{1,2,...,T-1\}$ (4)
@@
The reverse diffusion process gradually denoises standard Gaussian noise into the original data distribution by
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))+\sqrt{\beta_{t}}\epsilon,$ (6)
where $x_{T}\sim\mathcal{N}(0,I)$. After $T$ denoise steps, DDPM can generate the samples from the learned data distribution.
@@ III-B Problem Formulation
The objective of the multi-agent trajectory prediction task is to use the observed history trajectories to jointly generate future trajectories of all the agents in the scene. We denote the accurate and complete history trajectory as $c^{i}=\{S_{-T_{his}+1}^{i},S_{-T_{his}+2}^{i},\dots,S_{0}^{i}\}$ where $S_{k}^{i}$ stands for the two-dimensional position coordinate of $i$th agent at time step $k$. We denote the future trajectory as $y^{i}=\{P_{1}^{i},\dots,P_{T_{fut}}^{i}\},i=1,\dots,N_{a}$. The task is to generate the future trajectory distribution $p(y|c)$. However, the historical trajectory may be missing and injected with noise caused by unexpected conditions such as sensor failure and poor weather. We use an elementwise mask to represent missing trajectory data. We model the injected noise as Gaussian noise so the noisy trajectory is $c+n$, where $n\sim\mathcal{N}(0,\sigma I)$. $\sigma$ is the scale of Gaussian noise.
@@ III-C Represent Full Trajectory Distribution with Diffusion
Previous diffusion-based work [13, 22] formulated it as the conditional prediction process that uses the history trajectory $c$ as the condition to generate the future trajectory distribution $p(y|c)$. Specifically, their denoise module learns to predict the conditional noise value $\epsilon(y_{t},t|c)$. Our key insight here is we can re-formulate the prediction task solving an inverse problem of inpainting, which aims to generate full-length trajectories containing both history and future parts while ensuring the history parts are assembled to the corresponding observed history trajectories. We denote the generated history trajectory as $z_{i}$ and the corresponding full-length trajectory to be $x^{i}=z^{i}\cup y^{i}=\{S_{-T_{his}}^{i},\dots,S_{0}^{i},P_{1}^{i},\dots,P_{T_{fut}}^{i}\}$. We train the DDPM by Eq. (5) where $\epsilon$ has the same shape of $x={x^{1},\dots,x^{N_{a}}}$. Hence we approached to model the unconditional distribution $p(x)$, the one that we can sample from through direct denose steps Eq. (6) at inference time. As the objective remains to be sampling from $p(y|c)$, we first consider adding the conditions by shaping the score-matching process with proper guidance so that we acquire sample from $p(x|c)$. Sequentially, we take the future parts from sampled trajectories by $y=\phi_{fut}(x)$.
@@ III-D Robust Prediction as Posterior Sampling
During inference time, diffusion models sample from the prior distribution by iteratively recovering data from noise, which is a Markovian process described by Eq. (6). For various problems, we have the condition $c$ and the goal is to sample from posterior $p(x|c)$. According to Bayes’ rule: $p(x|c)=p(c|x)p(x)/p(c)$, we have the score of the posterior distribution:
$\nabla_{x}{\log{p_{t}}(x|c)}=\nabla_{x}{\log{p_{t}}(x)}+\nabla_{x}{\log{p_{t}}(c|x)}.$ (7)
According to diffusion posterior sampling (DPS) [16], if the condition or measurement has the form of a general noisy inverse problem $y=\phi(x_{0})+n$, where $\phi$ is an arbitrary operator, and $n\sim\mathcal{N}(0,\sigma^{2}I)$ is the Gaussian noise, then $\nabla_{x}{\log{p_{t}}(c|x)}$ can be approximated as
$\nabla_{x}{\log{p_{t}}(y|x)}=-\lambda\nabla_{x_{t}}\left\|y-\phi(\hat{x_{0}}(x_{t}))\right\|_{2}^{2}=-\lambda\nabla_{x_{t}}\mathcal{L},$ (8)
@@
$\hat{x_{0}}(x_{t})=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(x_{t}+(1-\overline{\alpha}_{t})\epsilon_{\theta}(x_{t},t)).$ (9)
Once the condition operator $\phi$ is known (i.e., $\phi_{his}(x)$ denotes selecting history dimensions from $x$), we can inject such condition by iteratively adding the guidance term Eq. (8) to the intermediate noisy data during the reverse diffusion process so that we drag the sample to the direction that minimizes the guidance loss $\mathcal{L}$. We refer to this characteristic as ”soft-conditioning”, which does not guarantee the exact inpainting of history trajectories. We will show such a design greatly boosts robustness and adaptation abilities in the following sections.
For conventional prediction tasks where we expect data to be clean and complete, soft-conditioning however, may result in ill-conditioned prediction. To further enhance performance on these tasks, we can strictly enforce history conditions following RePaint[24]. This plug-and-play modification introduces an additional step that first compromises observed history trajectories to the noise level $t$ through the forward diffusion process Eq. (4), then concatenates them with the corresponding future parts of intermediate samples.
@@
3:     $\epsilon\sim\mathcal{N}(0,\Sigma)$
4:     $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))+\sqrt{\beta_{t}}\epsilon$
5:     $\hat{x}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t))$
@@
9:        $\epsilon^{\prime}\sim\mathcal{N}(0,\Sigma)$

### v2

@@ Abstract
Joint pedestrian trajectory prediction has long grappled with the inherent unpredictability of human behaviors. Recent works employing conditional diffusion models in trajectory prediction have exhibited notable success. Nevertheless, the heavy dependence on accurate historical data results in their vulnerability to noise disturbances and data incompleteness. To improve the robustness and reliability, we introduce the Guided Full Trajectory Diffuser (GFTD), a novel diffusion-based framework that translates prediction as the inverse problem of spatial-temporal inpainting and models the full joint trajectory distribution which includes both history and the future. By learning from the full trajectory and leveraging flexible posterior sampling methods, GFTD can produce accurate predictions while improving the robustness that can generalize to scenarios with noise perturbation or incomplete historical data. Moreover, the pre-trained model enables controllable generation without an additional training budget. Through rigorous experimental evaluation, GFTD exhibits superior performance in joint trajectory prediction with different data quality and in controllable generation tasks. See more results at https://sites.google.com/andrew.cmu.edu/posterior-sampling-prediction.
@@ I INTRODUCTION
Pedestrian trajectory prediction is crucial for human-robot interaction systems such as autonomous driving, etc. The goal is to predict future trajectories based on previous pedestrian movements and environmental contexts. By accurately predicting pedestrian trajectories, autonomous systems can plan their actions accordingly, ensuring safe and efficient navigation in dynamic environments. However, the unpredictable and complicated nature of human behaviors makes this task challenging, especially with multiple agents involved where their interactions need to be considered.
@@
To simplify the problem, early works [1] [2] [3] [4] focused on marginal pedestrian trajectory prediction, which forecasts the trajectory for each pedestrian independently. Such approaches require a downstream planning module to perform safety checks for every combination of the individual predictions. Even so, combination rollouts could still produce unrealistic self-collisions and lead to failure in challenging scenarios. As a result, joint pedestrian trajectory prediction, which predicts consistent trajectories for all agents together, has gained attention in the community. [5] introduced a joint metrics term into supervision loss, transforming the marginal predictor into a joint one. However, joint pedestrian trajectory prediction remains challenging since existing approaches heavily rely on accurate and complete historical data to incorporate temporal and social dynamics. This dependence results in their vulnerability to interference from noisy disturbances and incomplete data, significantly threatening their effectiveness in real-world applications, such as sensors under adverse weather conditions.
@@
Figure 1: Illustration of existing trajectory prediction framework and our guided full trajectory diffuser framework. (a) Multi-agent trajectory prediction methods directly generate entire future trajectories through supervised learning. (b) Diffusion-based Multi-agent trajectory prediction methods generate future trajectories step-by-step during the denoising process. (c) Our method for multi-agent trajectory prediction incorporates history guidance within the diffusion framework. It predicts entire trajectories and supports additional capabilities in a single model, including Noise-Robust Trajectory Prediction, Incomplete Data Prediction, and Controllable Trajectory Generation.
@@
To deal with the noises, previous research involved augmenting data with predefined noise and training models on noisy datasets [6], and it can be further improved by adversarial training procedures [7]. In addition, historical trajectories from sensors could be incomplete. Previous studies  [8] [9] [10] [11] proposed to reconstruct incomplete data and predict future trajectories during the training phase, requiring a specifically designed model and well-established training strategy. Such approaches optimized per-problem functions, lacking adaptability to diverse tasks across various contexts.
@@
Drawing inspiration from the diffusion model with its remarkable capability of capturing the complicated distribution, we propose a unified framework, named, Guided Full Trajectory Diffuser (GFTD) for the joint pedestrian trajectory prediction to better handle the disturbances and incompleteness. GFTD represents the entire trajectory distribution, both historical and future, with one diffusion model. We formulate trajectory prediction and controllable generation as inverse problems and solve them through posterior sampling techniques. Specifically, we sample in-distribution full trajectories based on historical trajectories and priors such as physical constraints and behavioral intentions. Without the necessity for explicit training to handle noisy and/or incomplete inputs, GFTD enables robust prediction and controllable generation—both achievable during the inference time. In a nutshell, our proposed framework streamlines the training process and offers adaptability to various scenarios at inference, providing a solution that can address all challenges without extra training requirements.
@@ II-A Pedestrian Trajectory Prediction
Pedestrian trajectory prediction is crucial for many downstream tasks in autonomous driving, such as tracking [12] and mapping [13]. This task involves forecasting the future movement paths of pedestrians, given their historical movements and the environment. However, it is challenging to predict their motion because of the diverse and unpredictable nature of human behaviors. To address this issue, two mainstream paradigms have been developed. The supervised learning approach [4] aims to minimize the differences between the ground truth trajectories and the predictions using L2 loss, etc. On the other hand, the generative learning approach [1] formulates the prediction as a task of generating conditional distributions of the future trajectory based on past trajectories. Among various generative models, the diffusion model [14] [15] has shown exceptional performance in pedestrian trajectory prediction. Through a conditional reverse diffusion process, the diffusion model generates future trajectory distributions from a standard Gaussian distribution, which captures accurate and diverse predictions of future trajectories.
II-B Joint Trajectory Distribution Modelling
Recently, joint pedestrian trajectory prediction has gained significant attention. Unlike marginal trajectory prediction, which treats each pedestrian independently and can result in self-colliding trajectories between agents, joint trajectory prediction considers the interactions between future trajectories of agents, leading to more consistent and feasible predictions. [5] incorporates joint metrics as the training objective, transforming the marginal trajectory predictor into a joint trajectory predictor. In this paper, we propose Guided Full Trajectory Diffuser (GTFD) for joint pedestrian trajectory prediction, leveraging the strengths of diffusion models in generating accurate and diverse future trajectory distributions while considering the interactions between pedestrians.
@@ III Methods
In this section, we undertake a comprehensive exposition of our novel multi-agent trajectory prediction framework, conceptualizing the prediction task through spatial-temporal inpainting. We first introduce preliminaries on diffusion models and problem definition. Subsequently, we explain how we formulate our Guided Full Trajectory Diffuser (GFTD).
@@ III-A Preliminaries on Diffusion models
According to Itô Stochastic Differential Equations (SDE) [24], the forward data noising process is defined as the following form:
$dx=f(x,t)dt+g(t)dw.$ (1)
where $x$ represents the data, $w$ is the standard Wiener process, $f$ is the drift coefficient, $g$ is the diffusion coefficient, $t\in[0,T]$ is the diffusion step.
@@
$dx=[f(x,t)-g^{2}(t)\nabla_{x}{\log{p_{t}}(x)}]dt+g(t)dw.$ (2)
where $p_{t}(x)$ is the probability density, and $\nabla_{x}{\log{p_{t}}(x)}$ is the score function, which can be learned by a neural network $s_{\theta}(x_{t},t)$ with score matching. Here, $x_{t}$ denotes the intermediate noisy data $x$ at denoise time $t$. The training objective is defined as
$\mathbb{E}_{t}\{\lambda(t)\mathbb{E}_{x_{0}}\mathbb{E}_{x_{t}|x_{0}}[\left\|s_{\theta}(x_{t},t)-\nabla_{x_{t}}{\log{p}(x_{t}|x_{0})}\right\|_{2}^{2}]\}.$ (3)
where $\lambda(t)$ is a weighting function.
In this paper, we follow the implementation introduced by Denoising Diffusion Probabilistic Models (DDPM) [25]. DDPM gradually add noise to the original data $x$ according to a variance schedule $\beta_{1}$, $\beta_{2}$,…,$\beta_{T}$:
$x_{t}=\sqrt{1-\beta_{t}}x_{t-1}+\sqrt{\beta}_{t}\epsilon,\quad t\in\{1,2,...,T\}$ (4)
@@
With the learned $\epsilon_{\theta}(x_{t},t)$, the reverse diffusion process gradually denoises standard Gaussian noise into the original data distribution by
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))+\sqrt{\beta_{t}}\epsilon.$ (6)
where $x_{T}\sim\mathcal{N}(0,I)$, $t$ from $T$ to $1$.
@@ III-B Problem Formulation
The objective of the multi-agent trajectory prediction task is to use the observed history trajectories to jointly predict future trajectories of all the agents in the scene. We denote $P_{k}^{i}$ as the positions of agent $i$ at time step $k$, $i=1,\dots,N_{a}$. Joint history trajectory is $c=\{c^{i}\}_{i=1}^{N_{a}}$, $c^{i}=\{P_{-T_{his}+1}^{i},P_{-T_{his}+2}^{i},\dots\,P_{-1}^{i}\}$. Joint current position is $E=\{E^{i}\}_{i=1}^{N_{a}}$, $E^{i}=P_{0}^{i}$. JToint future trajectory is $y=\{y^{i}\}_{i=1}^{N_{a}}$, $y^{i}=\{P_{1}^{i},\dots,P_{T_{fut}}^{i}\}$. The task is to generate the future trajectory distribution $p(y|E,c)$. However, the historical trajectory may be incomplete and injected with noise caused by unexpected conditions such as sensor failure and poor weather. We use an elementwise mask to represent incomplete trajectory data and model the injected noise as Gaussian noise.
@@ III-C Represent Full Trajectory Distribution with Diffusion
Previous diffusion-based work [14] [23] [26] formulated prediction task as the diffusion process that conditioned on history trajectories $c$ and current position $E$ to generate future trajectory distribution $p(y|E,c)$. Our key insight here is that we can learn to generate full-length trajectories containing both the past and the future and regard the prediction task as an inverse problem of inpainting, which aims to infer and complete the incomplete trajectory (e.g., the future) based on observations (e.g., the past). We denote the generated history trajectory as $\hat{c}_{i}$ and the corresponding full-length trajectory as $x^{i}=\hat{c}^{i}\cup y^{i}=\{P_{-T_{his}}^{i},\dots,P_{-1}^{i},P_{1}^{i},\dots,P_{T_{fut}}^{i}\}$. We train the DDPM by Eq. (5) where $x=\{x^{i}\}_{i=1}^{N_{a}}$, and we can generate full trajectory distribution $p(x|E)$ by learned $\epsilon_{\theta}(x_{t},t,E)$ through Eq. (6).
@@ III-D Robust Prediction as Posterior Sampling
Prediction task is to obtain $p(y|E,c)$ which is equivalent to $p(y,c|E,c)=p(x|E,c)$. Given learned full trajectory $p(x|E)$ through diffusion model and Bayes’ rule, we have $p(x|E,c)=p(c|x,E)p(x|E)/p(c)$. Thus, the score of the posterior distribution can be calculated as:
$\nabla_{x}{\log{p_{t}}(x|E,c)}=\nabla_{x}{\log{p_{t}}(x|E)}+\nabla_{x}{\log{p_{t}}(c|x,E)}.$ (7)
According to diffusion posterior sampling (DPS) [17], if the condition or measurement has the form of a general noisy inverse problem $c=\phi(x_{0})+n$, where $\phi$ is an arbitrary operator, and $n\sim\mathcal{N}(0,\sigma^{2}I)$ is the Gaussian noise, then $\nabla_{x}{\log{p_{t}}(c|x,E)}$ can be approximated as
$\nabla_{x}{\log{p_{t}}(c|x,E)}=-\lambda\nabla_{x_{t}}\left\|c-\phi(\hat{x_{0}}(x_{t}))\right\|_{2}^{2}=-\lambda\nabla_{x_{t}}\mathcal{L},$ (8)
@@
$\hat{x_{0}}(x_{t})=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(x_{t}+(1-\overline{\alpha}_{t})\epsilon_{\theta}(x_{t},t,E)).$ (9)
Once the condition operator $\phi$ is known (i.e., $\phi_{his}(x)$ denotes historical portion of $x$), we can inject such condition by iteratively adding the guidance term Eq. (8) to the intermediate noisy data during the reverse diffusion process so that we drag the sample to the direction that minimizes the guidance loss $\mathcal{L}$ that reflects our preferences, which would be having the generated history trajectories close to the ground truth observation. Specific designs of $\mathcal{L}$ will be elaborated in section IV. With the guidance gradient restricted to a certain range, the generated noisy samples will be lying in the data manifolds, thus avoiding generating out-of-distribution samples. Even though this does not guarantee the exact recovery of history trajectories and may result in some degrees of ill-conditioned prediction, a properly guided sample still achieves high prediction accuracy. We refer to this characteristic as ”soft-conditioning” and will show such a design greatly boosts robustness and adaptation abilities in the following sections.
For conventional prediction tasks, we expect data to be clean and complete and do not need to trade condition correctness off for perturbation tolerance. To further enhance performance on these tasks, we can strictly enforce history conditions following RePaint[27]. This plug-and-play modification can be manually enabled. It introduces an additional step that first compromises observed history trajectories to the noise level $t$ through the forward diffusion process Eq. (4), then concatenates them with the corresponding future parts of intermediate samples.
@@
3:     $\epsilon\sim\mathcal{N}(0,I)$
4:     $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t,E))+\sqrt{\beta_{t}}\epsilon$
5:     $\hat{x}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t,E))$
@@
9:        $\epsilon^{\prime}\sim\mathcal{N}(0,I)$